OpenStudy (anonymous):

Solve: ) open parentheses 6 plus square root of 2 close parentheses open parentheses 6 minus square root of 2

hero (hero):

Hint: \[(6 + \sqrt{2})(6 - \sqrt{2})\]

hero (hero):

2nd Hint: (a + b)(a - b) = a^2 - b^2

OpenStudy (anonymous):

Is it 36 - the square root of 2?

hero (hero):

Correction: 36 - 2

OpenStudy (anonymous):

the answer is 34

hero (hero):

or simply 34